A 75-g ice cube at 0°C is heated until 68.4 g has become water at 100°C and 6.6 g has become steam at 100°C. How much energy was added to accomplish the transformation? kJ
Answer #1


heat added = heat required to melt 68.4 g of ice + heat required to convert 6.6 g to steam

heat required for 68.4 of ice to melt to water at 100 Q1 = (Mw*L)+(Mw*Sw*dT)

Lf = 334 J/g

Sw = 4.186 J/g

Lv = 2230 J/g

Q1 = (68.4*334) + (68.4*4.186*100)

Q1 = 51477.84 J


heat required to convert 6.6 g to steam Q2 = Msteam*( Lf + (Sw*dT) + Lv )

Q2 = 6.6*( 334 + (4.186*100) + 2230 )

Q2 = 19685.16 J

Q = Q1 +Q2

Q = 51477.84 + 19685.16 = 71163 J

Q = 71.1 KJ